Intro Graphics February 13, 2006 G. Farin

Programming Assignment # 2 due: 2-27 percent of your your grade: 12.5

Create a 3D viewer for an object, represented as a triangle mesh. You should be able to rotate the object using the mouse, using perspective or orthographic projection. Have a function for toggling face normals off or on. Have a function for displaying wire frame or flat shaded. Have a function for zooming in or out.

Work with you own example file to start with – you will be given a ”real” object later.